The difference between simple compound interest on rs 1625 for 3 year at 4% per annum in rupees is?
Answer by fractalier(6550)

Compound interest - simple interest =
P(1+r)^t - (P+Prt) =
1625(1.04)^3 - (1625+1625(.04)(3))=
1827.90 - 1820 = 7.90 rupees